# Content-Oblivious Leader Election in 2-Edge-Connected Networks

### v2

@@ 4.2 DFS Notification
We now proceed to analyze the DFS Notification phase. This phase of the algorithm employs a DFS to notify all nodes of the elected leader, while guaranteeing termination with quiescence. Although a graph can have many DFS trees, we use the specific tree $T_{G}$, rooted at the node with the smallest identifier and prioritizing smaller port numbers during traversal, see Section 2.3. To begin, observe that a DFS visits each edge of the graph exactly once in each direction, as detailed below.
@@
We now discuss how the DFS traversal is implemented in Algorithm 1 and Algorithm 2.
@@
The use of $\mathsf{SendPulsesUntil}_{j}(\mathsf{Leader}\mathsf{ID}+N+2)$ for sending both types of notifications serves to ensure quiescence upon termination, as it allows each node to anticipate and recognize the precise number of pulses—$\mathsf{Leader}\mathsf{ID}+N+2$—it will receive before becoming inactive.
@@
$\mathcal{E}_{u\rightarrow v}^{\mathsf{Send}}$ =$=\begin{cases}\mathsf{SendExplore}_{i}(u)&\text{if $(u,v)$ is before $(v,u)$ in $A$,}\\ \mathsf{SendDone}_{i}(u)&\text{otherwise.}\end{cases}$
$\mathcal{E}_{u\rightarrow v}^{\mathsf{Receive}}$ =$=\begin{cases}\mathsf{ReceiveExplore}_{j}(v)&\text{if $(u,v)$ is before $(v,u)$ in $A$,}\\ \mathsf{ReceiveDone}_{j}(v)&\text{otherwise.}\end{cases}$
$t_{u\rightarrow v}^{\mathsf{Send}}$ $=\text{the time when the event $\mathcal{E}_{u\rightarrow v}^{\mathsf{Send}}$ happens.}$.
$t_{u\rightarrow v}^{\mathsf{Receive}}$ $=\text{the time when the event $\mathcal{E}_{u\rightarrow v}^{\mathsf{Receive}}$ happens.}$.
$t_{r}^{\mathsf{StartDFS}}$ $=\text{the time when the event $\mathsf{StartDFS}(r)$ happens.}$.
@@
Recall that $P_{v,r}$ is a shortest path from $v$ to $r$ in $\vec{G}$, see Section 2.3.
For the base case of $i=0$, $\mathsf{IH}(0)$ follows from Lemma 14. Specifically, for $i=0$, (H1) asserts that $t_{0}=t_{r}^{\mathsf{StartDFS}}<\infty$, (H2) asserts that all events in $\mathscr{E}^{\star}\setminus\left\{t_{r}^{\mathsf{StartDFS}}\right\}$ do not occur before $t_{0}=t_{r}^{\mathsf{StartDFS}}$, and (H3) asserts that $\mathsf{ID}(r)=\mathsf{Count}(r)$ at the time $t_{0}=t_{r}^{\mathsf{StartDFS}}$, since $R_{0}=\{r\}$.
@@
Observation 15 (Next sending occurs immediately).
@@
By (H1)$(i)$ and (H2)$(i)$, the execution of the algorithm follows the desired DFS traversal for the first $i$ directed edges, where the last event $\mathcal{E}$ is $\mathsf{StartDFS}(r)$ if $i=0$ and is $\mathcal{E}_{u_{i}\rightarrow v_{i}}^{\mathsf{Receive}}$ if $i>0$. According to the description of our algorithm, immediately after triggering the last event $\mathcal{E}$, the current node $v_{i+1}$ is ready to continue the DFS traversal by passing the token along the next directed edge $e_{i+1}=(u_{i+1},v_{i+1})$, triggering $t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}$, so the first statement holds.
Other than $\mathcal{E}$ and $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}$, no other event in $\mathscr{E}^{\star}$ associated with the current node $v_{i+1}$ is triggered at time $t_{i}$. The local computation at $v_{i+1}$ cannot trigger any event in $\mathscr{E}^{\star}$ associated with other nodes. Hence the second statement holds. ∎
15 implies that, in the proof of the inductive step below, to establish (H1)$(i+1)$ assuming $\mathsf{IH}(i)$, we may assume that the event $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}$ has already occurred, and focus on proving that the corresponding event $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ eventually occurs.
@@
Observation 16 (Correct $\mathsf{Leader}\mathsf{ID}$).
@@
Observation 17 (DFS notification pulses).
@@
We discuss one subtle issue in the proof: While the upper bound $\mathsf{ID}(r)+N+2$ appears to be clear, there might be a possibility that $u$ already sends to $v$ more than this number of pulses while $u$ is in the synchronized counting phase. This is however impossible because the number of such pulses is $\mathsf{Count}(u)+1\leq(\mathsf{ID}(r)+N-1)+1<\mathsf{ID}(r)+N+2$ according to 9 and (H3)$(i)$. This also implies that some of the $\mathsf{ID}(r)+N+2$ pulses sent from $u$ to $v$ must be DFS notification pulses. ∎
@@
Before proving Lemma 18, we first use it to prove the main theorem.
Proof of Theorem 1.
The base case $\mathsf{IH}(0)$ follows from Lemma 14, which shows that $\mathsf{StartDFS}(r)$ occurs after finite time and is before all other events in $\mathscr{E}^{\star}$, satisfying (H1) and (H2). The condition (H3) is satisfied because $r$ enters the DFS notification phase through the $\mathsf{Leader}$ exit condition, which ensures $\mathsf{Count}(r)=\mathsf{ID}(r)=\mathsf{ID}(r)+P_{r,r}$, as $P_{r,r}=0$.
By Lemma 18, $\mathsf{IH}(1),\mathsf{IH}(2),\ldots,\mathsf{IH}(2|E|)$ hold. Given that $\mathsf{IH}(2|E|)$ holds, we show that the algorithm terminates by time $t_{2|E|}$ with quiescence correctly. By 16, all nodes $v$ have advanced to the DFS notification phase with $\mathsf{Leader}\mathsf{ID}(v)=\mathsf{ID}(r)\neq\mathsf{ID}(v)$. This also implies that all of the nodes in $V\setminus\{r\}$ are non-leaders, so $r$ is the unique leader.
Given that $\mathsf{IH}(2|E|)$ holds, each edge is traversed exactly once in both directions during the DFS. For each node $v$, $\mathcal{E}_{u\rightarrow v}^{\mathsf{Receive}}$ has been triggered for each of its neighbors $u$, meaning that $\mathcal{P}(v)=\emptyset$ and $v$ has terminated by the time $t_{2|E|}$. Moreover, the termination comes with a quiescence guarantee by 16.
The message complexity bound follows from 16, which states that the number of pulses sent along each direction of each edge is exactly $\mathsf{ID}(r)+N+2$, where $\mathsf{ID}(r)=\mathsf{ID_{\min}}\cdot N$, as we multiply each node identifier by $N$ at the beginning. Therefore, the overall message complexity is $O(m\cdot N\cdot\mathsf{ID_{\min}})$.
@@
For the rest of the discussion, we prove Lemma 18. Suppose $\mathsf{IH}(i)$ holds. Our goal is to show that $\mathsf{IH}(i+1)$ holds. By 15, we may assume that
@@
and take it as the starting point of the analysis. We want to show that $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ happens eventually, before triggering any other events in $\mathscr{E}^{\star}$. The proof consists of two parts. In the first part, we show that, before $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$, no other event in $\mathscr{E}^{\star}$ can happen, so (H2)$(i+1)$ holds. We also show that the following small counter property holds for not only the nodes in the DFS notification phase but also the remaining nodes in the synchronized counting phase. In the second part, we show that $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ occurs within finite time, so (H1)$(i+1)$ holds. Moreover, if this is also the moment where $v_{i+1}$ transitions from the synchronized counting phase to the DFS notification phase, then the counter value of $v_{i+1}$ satisfies H3$(i+1)$.
@@
Suppose $\mathsf{IH}(i)$ holds. Consider a moment $t\geq t_{i}$. Suppose $\mathcal{E}_{B}$ has not happened, then at any moment within the time interval $[t_{i},t]$, there is no DFS notification pulse in transit, except for the directed edge $e_{i+1}=(u_{i+1},v_{i+1})$.
@@
By (H3)$(i)$, $\mathsf{Count}(w)\leq\mathsf{ID}(r)+\mathsf{length}(P_{w,r})$. By Lemma 11, $\mathsf{Count}(v)$ is at most
@@
Suppose $\mathsf{IH}(i)$ holds. Consider a moment $t\geq t_{i}$. If both $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ have not happened and the following statements hold.
@@
By Lemma 20, the next pulse delivered—as well as any pulse delivered after time $t_{i}$ and before time $t$—is either a DFS notification pulse along the directed edge $e_{i+1}=(u_{i+1},v_{i+1})$ or a synchronized counting pulse. Combining this with 17, the precondition of Lemma 21 is satisfied at the moment after the next pulse delivery, so Lemma 21 implies that the next pulse delivery does not trigger $\mathcal{E}_{A}$.
@@
If the pulse is a DFS notification pulse along the directed edge $e_{i+1}=(u_{i+1},v_{i+1})$, then $\mathcal{E}$ must be $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$. Therefore, for the rest of the discussion, we assume that the pulse is a synchronized counting pulse. In particular, Lemma 12 implies that the pulse cannot trigger the $\mathsf{Non\text{-}Leader}$ exit condition.
Therefore, the only remaining way to trigger any event in $\mathscr{E}^{\star}$ is to get a total of $\mathsf{ID}(r)+N+2$ pulses via some port, but that is not possible with synchronized counting pulses only, unless the small counter property is violated. Specifically, 9 implies that the number of synchronized counting pulses sent along a port equals the counter value plus one. While the small counter property holds, this number if capped at $(\mathsf{ID}(r)+N-1)+1$, which is smaller than the required $\mathsf{ID}(r)+N+2$.
@@
For the base case of the induction, consider the moment $t=t_{i}$. By 15, $\mathcal{E}_{B}$ has not happened yet. By a combination of 17 and Lemma 20, the precondition of Lemma 21 is satisfied at $t=t_{i}$, so $\mathcal{E}_{A}$ has not happened yet.
For the inductive step, apply Lemma 22 to any moment $t\in\left[t_{i},t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}\right)$. Suppose, by induction hypothesis that both $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ have not happened by time $t$, then it is guaranteed that the next pulse delivery does not trigger $\mathcal{E}_{A}$. Moreover, if the next pulse delivery does not trigger $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$, then it also does not trigger $\mathcal{E}_{B}$. Therefore, $\mathcal{E}_{A}$ does not occur in the time interval $\left[t_{i},t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}\right]$, and $\mathcal{E}_{B}$ does not occur in the time interval $\left[t_{i},t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}\right)$. ∎
@@
Let the port of the edge $\{u_{i+1},v_{i+1}\}$ be $j$ in $u_{i+1}$ and $k$ in $v_{i+1}$. By 15, $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}$ occurs at time $t_{i}$. We break the analysis into a few cases.
@@
Suppose $\{u_{i+1},v_{i+1}\}$ is a back edge, meaning that $v_{i+1}$ has already advanced to the DFS notification phase by time $t_{i}$. By (H3)$(i)$, $\mathsf{Count}(v_{i+1})\geq\mathsf{ID}(r)$, so $v_{i+1}$ has already sent to $u_{i+1}$ at least $\mathsf{Count}(v_{i+1})+1\geq\mathsf{ID}(r)+1$ synchronized counting pulses (9), meaning that $u_{i+1}$ only waits for a finite amount of time, and eventually $v_{i+1}$ will receive a total of $\mathsf{ID}(r)+N+2$ pulses from $u_{i+1}$, triggering $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$.
@@
First, we claim that the waiting before sending pulses takes finite time for $u_{i+1}$ by applying Lemma 13 to the moment $t_{i}$ with $S=V\setminus R_{i}$. At least one of the three types of progress is made within finite time.
@@
    where the first inequality follows from (H3)$(i)$. By 9, this means that $v_{i+1}$ has sent out $\mathsf{ID}(r)+1$ pulses to $u_{i+1}$, and eventually these pulses will arrive at $u_{i+1}$, so $u_{i+1}$ only waits for a finite time.
@@
    is triggered. In the former case, we are done already. The latter case is not possible before first triggering $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ in view of Lemma 23.
  • Suppose (C3) holds, meaning that a DFS notification pulse from some node in $V\setminus S$ is delivered to some node in $S$. By a combination of Lemma 20 and Lemma 23, we infer that unless $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ has been triggered, the pulse must be sent from $u_{i+1}$ to $v_{i+1}$. In either case, the waiting must be done within finite time.
@@
For $v_{i+1}$ to confirm the receipt of the explore-notification, we need $v_{i+1}$ to first satisfy the $\mathsf{Non\text{-}Leader}$ exit condition, To do so, we use the small counter property, which is guaranteed to hold during the time interval $\left[t_{i},t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}\right)$ in view of Lemma 23.
Due to the small counter property, the counter value of $v_{i+1}$ is capped by $\mathsf{ID}(r)+N-1$, so by 9, while $v_{i+1}$ is still in the synchronized counting phase,
@@
Now we have collected all the needed ingredients to prove Lemma 18.
Proof of Lemma 18..
@@
The part $t_{i}=t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}<t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ follows from 15. The part $t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}<\infty$ follows from Lemma 24. We set $t_{i+1}=t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ To prove (H2)$(i+1)$, it suffices to show that, within the timer interval $[t_{i},t_{i+1})$, all events in $\mathcal{E}_{B}$ do not occur. This follows from Lemma 23. To prove (H3)$(i+1)$, it suffices to show that

### v3

@@ 4.2 DFS Notification
We now proceed to analyze the DFS Notification phase. This phase of the algorithm employs a DFS to notify all nodes of the elected leader, while guaranteeing termination with quiescence. Although a graph can have many DFS trees, we use the specific tree $T_{G}$, rooted at the node with the smallest identifier and prioritizing smaller port numbers during traversal, see Section˜2.3. To begin, observe that a DFS visits each edge of the graph exactly once in each direction, as detailed below.
@@
We now discuss how the DFS traversal is implemented in Algorithm˜1 and Algorithm˜2.
@@
[…]
@@
$\mathcal{E}_{u\rightarrow v}^{\mathsf{Send}}$ $=\begin{cases}\mathsf{SendExplore}_{i}(u)&\text{if $(u,v)$ is before $(v,u)$ in $A$,}\\ \mathsf{SendDone}_{i}(u)&\text{otherwise.}\end{cases}$
$\mathcal{E}_{u\rightarrow v}^{\mathsf{Receive}}$ $=\begin{cases}\mathsf{ReceiveExplore}_{j}(v)&\text{if $(u,v)$ is before $(v,u)$ in $A$,}\\ \mathsf{ReceiveDone}_{j}(v)&\text{otherwise.}\end{cases}$
$t_{u\rightarrow v}^{\mathsf{Send}}$ $=\text{the time when the event $\mathcal{E}_{u\rightarrow v}^{\mathsf{Send}}$ happens.}$
$t_{u\rightarrow v}^{\mathsf{Receive}}$ $=\text{the time when the event $\mathcal{E}_{u\rightarrow v}^{\mathsf{Receive}}$ happens.}$
$t_{r}^{\mathsf{StartDFS}}$ $=\text{the time when the event $\mathsf{StartDFS}(r)$ happens.}$
@@
Recall that $P_{v,r}$ is a shortest path from $v$ to $r$ in $\vec{G}$, see Section˜2.3.
For the base case of $i=0$, $\mathsf{IH}(0)$ follows from Lemma˜15. Specifically, for $i=0$, (H1) asserts that $t_{0}=t_{r}^{\mathsf{StartDFS}}<\infty$, (H2) asserts that all events in $\mathscr{E}^{\star}\setminus\left\{t_{r}^{\mathsf{StartDFS}}\right\}$ do not occur before $t_{0}=t_{r}^{\mathsf{StartDFS}}$, and (H3) asserts that $\mathsf{ID}(r)=\mathsf{Count}(r)$ at the time $t_{0}=t_{r}^{\mathsf{StartDFS}}$, since $R_{0}=\{r\}$.
@@
Observation 16 (Next sending occurs immediately).
@@
By (H1)$(i)$ and (H2)$(i)$, the execution of the algorithm follows the desired DFS traversal for the first $i$ directed edges, where the last event $\mathcal{E}$ is $\mathsf{StartDFS}(r)$ if $i=0$ and is $\mathcal{E}_{u_{i}\rightarrow v_{i}}^{\mathsf{Receive}}$ if $i>0$. According to the description of our algorithm, immediately after triggering the last event $\mathcal{E}$, the current node $v_{i}=u_{i+1}$ is ready to continue the DFS traversal by passing the token along the next directed edge $e_{i+1}=(u_{i+1},v_{i+1})$, triggering $t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}$, so the first statement holds.
Other than $\mathcal{E}$ and $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}$, no other event in $\mathscr{E}^{\star}$ associated with the current node $v_{i}=u_{i+1}$ is triggered at time $t_{i}$. The local computation at $v_{i}=u_{i+1}$ cannot trigger any event in $\mathscr{E}^{\star}$ associated with other nodes. Hence the second statement holds. ∎
˜16 implies that, in the proof of the inductive step below, to establish (H1)$(i+1)$ assuming $\mathsf{IH}(i)$, we may assume that the event $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}$ has already occurred, and focus on proving that the corresponding event $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ eventually occurs.
@@
Observation 17 (Correct $\mathsf{Leader}\mathsf{ID}$).
@@
Observation 18 (DFS notification pulses).
@@
We discuss one subtle issue in the proof: While the upper bound $\mathsf{ID}(r)+N+2$ appears to be clear, there might be a possibility that $u$ already sends to $v$ more than this number of pulses while $u$ is in the synchronized counting phase. This is however impossible because the number of such pulses is $\mathsf{Count}(u)+1\leq(\mathsf{ID}(r)+N-1)+1<\mathsf{ID}(r)+N+2$ according to ˜10 and (H3)$(i)$. This also implies that some of the $\mathsf{ID}(r)+N+2$ pulses sent from $u$ to $v$ must be DFS notification pulses. ∎
@@
Before proving Lemma˜19, we first use it to prove the main theorem.
Proof of Theorem˜1.
The base case $\mathsf{IH}(0)$ follows from Lemma˜15, which shows that $\mathsf{StartDFS}(r)$ occurs after finite time and is before all other events in $\mathscr{E}^{\star}$, satisfying (H1) and (H2). The condition (H3) is satisfied because $r$ enters the DFS notification phase through the $\mathsf{Leader}$ exit condition, which ensures $\mathsf{Count}(r)=\mathsf{ID}(r)=\mathsf{ID}(r)+P_{r,r}$, as $P_{r,r}=0$.
By Lemma˜19, $\mathsf{IH}(1),\mathsf{IH}(2),\ldots,\mathsf{IH}(2|E|)$ hold. Given that $\mathsf{IH}(2|E|)$ holds, we show that the algorithm terminates by time $t_{2|E|}$ with quiescence correctly. By ˜17, all nodes $v$ have advanced to the DFS notification phase with $\mathsf{Leader}\mathsf{ID}(v)=\mathsf{ID}(r)\neq\mathsf{ID}(v)$. This also implies that all of the nodes in $V\setminus\{r\}$ are non-leaders, so $r$ is the unique leader.
Given that $\mathsf{IH}(2|E|)$ holds, each edge is traversed exactly once in both directions during the DFS. For each node $v$, $\mathcal{E}_{u\rightarrow v}^{\mathsf{Receive}}$ has been triggered for each of its neighbors $u$, meaning that $\mathcal{P}(v)=\emptyset$ and $v$ has terminated by the time $t_{2|E|}$. Moreover, the termination comes with a quiescence guarantee by ˜17.
The message complexity bound follows from ˜17, which states that the number of pulses sent along each direction of each edge is exactly $\mathsf{ID}(r)+N+2$, where $\mathsf{ID}(r)=\mathsf{ID_{\min}}\cdot N$, as we multiply each node identifier by $N$ at the beginning. Therefore, the overall message complexity is $O(m\cdot N\cdot\mathsf{ID_{\min}})$.
@@
For the rest of the discussion, we prove Lemma˜19. Suppose $\mathsf{IH}(i)$ holds. Our goal is to show that $\mathsf{IH}(i+1)$ holds. By ˜16, we may assume that
@@
and take it as the starting point of the analysis. We want to show that $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ happens eventually, before triggering any other events in $\mathscr{E}^{\star}$. The proof consists of two parts. In the first part, we show that, before $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$, no other event in $\mathscr{E}^{\star}$ can happen, so (H2)$(i+1)$ holds. We also show that the following small counter property holds for not only the nodes in the DFS notification phase but also the remaining nodes in the synchronized counting phase. In the second part, we show that $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ occurs within finite time, so (H1)$(i+1)$ holds. Moreover, if this is also the moment where $v_{i+1}$ transitions from the synchronized counting phase to the DFS notification phase, then the counter value of $v_{i+1}$ satisfies H3(i+1)(i+1).
@@
Suppose $\mathsf{IH}(i)$ holds. Consider a moment $t\geq t_{i}$. Suppose $\mathcal{E}_{B}$ has not happened, then at any moment within the time interval $t_{i},t]$, there is no DFS notification pulse in transit, except for the directed edge $e_{i+1}=(u_{i+1},v_{i+1})$.
@@
By (H3)$(i)$, $\mathsf{Count}(w)\leq\mathsf{ID}(r)+\mathsf{length}(P_{w,r})$. By Lemma˜12, $\mathsf{Count}(v)$ is at most
@@
Suppose $\mathsf{IH}(i)$ holds. Consider a moment $t\geq t_{i}$. If both $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ have not happened, then the following statements hold.
@@
By Lemma˜21, the next pulse delivered—as well as any pulse delivered after time $t_{i}$ and before time tt—is either a DFS notification pulse along the directed edge $e_{i+1}=(u_{i+1},v_{i+1})$ or a synchronized counting pulse. Combining this with ˜18, the precondition of Lemma˜22 is satisfied at the moment after the next pulse delivery, so Lemma˜22 implies that the next pulse delivery does not trigger $\mathcal{E}_{A}$.
@@
If the pulse is a DFS notification pulse along the directed edge $e_{i+1}=(u_{i+1},v_{i+1})$, then $\mathcal{E}$ must be $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$. Therefore, for the rest of the discussion, we assume that the pulse is a synchronized counting pulse. In particular, Lemma˜13 implies that the pulse cannot trigger the $\mathsf{Non\text{-}Leader}$ exit condition.
Therefore, the only remaining way to trigger any event in $\mathscr{E}^{\star}$ is to get a total of $\mathsf{ID}(r)+N+2$ pulses via some port, but that is not possible with synchronized counting pulses only, unless the small counter property is violated. Specifically, ˜10 implies that the number of synchronized counting pulses sent along a port equals the counter value plus one. While the small counter property holds, this number if capped at $(\mathsf{ID}(r)+N-1)+1$, which is smaller than the required $\mathsf{ID}(r)+N+2$.
@@
For the base case of the induction, consider the moment $t=t_{i}$. By ˜16, $\mathcal{E}_{B}$ has not happened yet. By a combination of ˜18 and Lemma˜21, the precondition of Lemma˜22 is satisfied at $t=t_{i}$, so $\mathcal{E}_{A}$ has not happened yet.
For the inductive step, apply Lemma˜23 to any moment $t\in\left[t_{i},t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}\right)$. Suppose, by induction hypothesis that both $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$ have not happened by time $t$, then it is guaranteed that the next pulse delivery does not trigger $\mathcal{E}_{A}$. Moreover, if the next pulse delivery does not trigger $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$, then it also does not trigger $\mathcal{E}_{B}$. Therefore, $\mathcal{E}_{A}$ does not occur in the time interval $\left[t_{i},t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}\right]$, and $\mathcal{E}_{B}$ does not occur in the time interval $\left[t_{i},t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}\right)$. ∎
@@
Let the port of the edge $\{u_{i+1},v_{i+1}\}$ be $j$ in $u_{i+1}$ and $k$ in $v_{i+1}$. By ˜16, $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}$ occurs at time $t_{i}$. We break the analysis into a few cases.
@@
Suppose $\{u_{i+1},v_{i+1}\}$ is a back edge, meaning that $v_{i+1}$ has already advanced to the DFS notification phase by time $t_{i}$. By (H3)$(i)$, $\mathsf{Count}(v_{i+1})\geq\mathsf{ID}(r)$, so $v_{i+1}$ has already sent to $u_{i+1}$ at least $\mathsf{Count}(v_{i+1})+1\geq\mathsf{ID}(r)+1$ synchronized counting pulses (˜10), meaning that $u_{i+1}$ only waits for a finite amount of time, and eventually $v_{i+1}$ will receive a total of $\mathsf{ID}(r)+N+2$ pulses from $u_{i+1}$, triggering $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$.
@@
First, we claim that the waiting before sending pulses takes finite time for $u_{i+1}$ by applying Lemma˜14 to the moment $t_{i}$ with $S=V\setminus R_{i}$. At least one of the three types of progress is made within finite time.
@@
    where the first inequality follows from (H3)$(i)$. By ˜10, this means that $v_{i+1}$ has sent out $\mathsf{ID}(r)+1$ pulses to $u_{i+1}$, and eventually these pulses will arrive at $u_{i+1}$, so $u_{i+1}$ only waits for a finite time.
@@
    is triggered. In the former case, we are done already. The latter case is not possible before first triggering $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ in view of Lemma˜24.
  • Suppose (C3) holds, meaning that a DFS notification pulse from some node in $V\setminus S$ is delivered to some node in $S$. By a combination of Lemma˜21 and Lemma˜24, we infer that unless $\mathcal{E}_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ has been triggered, the pulse must be sent from $u_{i+1}$ to $v_{i+1}$. In either case, the waiting must be done within finite time.
@@
For $v_{i+1}$ to confirm the receipt of the explore-notification, we need $v_{i+1}$ to first satisfy the $\mathsf{Non\text{-}Leader}$ exit condition, To do so, we use the small counter property, which is guaranteed to hold during the time interval $\left[t_{i},t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}\right)$ in view of Lemma˜24.
Due to the small counter property, the counter value of $v_{i+1}$ is capped by $\mathsf{ID}(r)+N-1$, so by ˜10, while $v_{i+1}$ is still in the synchronized counting phase,
@@
Now we have collected all the needed ingredients to prove Lemma˜19.
Proof of Lemma˜19..
@@
The part $t_{i}=t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Send}}<t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ follows from ˜16. The part $t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}<\infty$ follows from Lemma˜25. We set $t_{i+1}=t_{u_{i+1}\rightarrow v_{i+1}}^{\mathsf{Receive}}$ To prove (H2)$(i+1)$, it suffices to show that, within the timer interval $t_{i},t_{i+1})$, all events in $\mathcal{E}_{B}$ do not occur. This follows from Lemma˜24. To prove (H3)$(i+1)$, it suffices to show that